Define FET parameters obtain relation between them
1
Expert's answer
2020-12-28T08:59:27-0500

Major FET parameter

  • Gate Cut Off Voltage (VGS(off))

The relation between gate to source voltage and drain current is given below.

"I_D = I_{DSS}[1-\frac{V_{GS}}{V_{GS(off)}}]^2"

  • Shorted Gate Drain Current ("I_{DSS}" )
  • Transconductance ("g_m" )

"g_m = \frac{\delta ID}{\delta V_{GS}}" at constant "V_{GS}"

  • Dynamic Output Resistance ("r_d" )

"g_m = \frac{\delta ID}{\delta V_{GS}}" at constant "I_D"

  • Amplification Factor (ΞΌ)

"\mu = \frac{\delta V_{DS}}{\delta V_{GS}}=\frac{\delta V_{DS}}{\delta I_D} \times \frac{\delta ID}{\delta V_{GS}}"

"\mu = r_d \times g_m"

Where,

"I_D =" drain current

"V_{GS} =" gate to source voltage